OpenStudy (anonymous):

Please help finish the sentence! When a complex number z is written in its polar form, z=r(cos theta + i sin theta), the non-negative number r is called ______ of z.

OpenStudy (anonymous):

If \(z=a+bi\) and \(r=\sqrt{a^2+b^2}\), then \(r=|z|\). Starts with an "m"...

OpenStudy (anonymous):

modulus?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yay! Thank you!

OpenStudy (anonymous):

You're welcome!